answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

When tossing 2 coins the probability of getting exactly 1 tail is one over 3?

No, the possible results of two coin flips are;

HH

HT

TH

TT

The probability of getting exactly one tail is 1/2 or 50%.

How many ways can you choose 2 books from 5 on a shelf?

You can wind up with 10 different pairs of books in your hand, which

you can choose from a shelf of 5 books in 20 different ways.

What are the names of the queens in a deck of cards?

Queen of Hearts, Queen of Diamonds, Queen of Spades, Queen of Clubs.

What is the or rule in probability?

Given two events, A and B, the probability of A or B is the probability of occurrence of only A, or only B or both.

In mathematical terms: Prob(A or B) = Prob(A) + Prob(B) - Prob(A and B).

What is the probability of three digit numbers without repeated digits with numbers 0123456?

It is possible to create a 3-digit number, without repeated digits so the probability is 1.

How many permutation of umbrella?

There are 20160 permutations of the word.

What is the probability of first rolling a 5 and then rolling a 3 using a fair six-sided die?

Probability of ' 5 ' on the first roll = 1/6

Probability of ' 3 ' on the second roll = 1/6

Probability of ' 5 ' followed by ' 3 ' = 1/6 x 1/6 = 1/36 = 27/9 percent

What percent of data falls between 1 Standard deviation below and 2 stand deviations above the mean?

The answer will depend on what the distribution is. Non-statisticians often assum that the variable that they are interested in follows the Standard Normal distribution. This assumption must be justified.

If that is the case then the answer is 81.9%

What is the probability of rolling and odd number on one toss of a number cube?

On a typical 6-sided number cube, there are three odd numbers (1, 3, 5) and three even numbers (2, 4, 6). Therefore, there are three outcomes that are desired out of a total of six possible outcomes. 3/6 = 1/2, so your probability is 1 in 2 or 50%.

What is the probability that exactly two times head are being tossed?

The answer depends on how many times the coin is tossed. The probability is zero if the coin is tossed only once!

Making some assumptions and rewording your question as "If I toss a fair coin twice, what is the probability it comes up heads both times" then the probability of it being heads on any given toss is 0.5, and the probability of it being heads on both tosses is 0.5 x 0.5 = 0.25.

If you toss it three times and want to know what the probability of it being heads exactly twice is, then the calculation is more complicated, but it comes out to 0.375.

How many ways can you rearrange small?

The letter is the word "small" can be rearranged in 60 different ways.

What is the probability that a couple will have 4 boys and 1 girl?

There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.

However, if you assume that they are independent events then, given that the probability of a boy is approx 0.52, the probability 4 boys and 1 girl out of 5 children is 0.1724 approx.

What is the probability of drawing a 6 or diamond from a deck of 52 cards?

The answer depends on how many cards are drawn and whether they are drawn at random.

The probability that a single card, drawn at random from a deck of cards in 16/52 = 4/13.

What is the difference between a simple random sample and a stratified random sample?

Sometimes a population consists of a number of subsets (strata) such that members within any particular strata are alike while difference between strata are more than simply random variations. In such a case, the population can be split up into strata. Then a stratified random sample consists of simple random samples, with the same sampling proportion, taken within each stratum.

Trending Questions
What is the chance of spinning a number greater than 1? If two dice are rolled Find the probability that the score on the dice is either 7 or 9? A spinner has the numbers 10 20 30 40 and a number cube numbered 1 through 6 what is the probability of getting a sum greater than 25? When 2 dice are rolled find the probability of getting a sum of 10 or doubles? A IRS auditor randomly selects 3 tax returns from 45 of which 15 contains errors What is the probability that she selects none of those containing errors? What does population size means? What is the proability of a factor of 12? On a math test 5 out of 20 students got an A If three students are chosen at random what is the probability that all three got an A on the test? What is the probability of choosing a letter from the word space? What is standard normal distribution of 0.62? What are some probability questions for a nine sided dice? What is the function of spade? What number do you have at the beginning when you multiply by 308 and point ten then divide by 3? There are 20 students participating in a math bee How many ways can the students be chosen to go first second third and fourth? What is the probability of picking a jack from a pack of playing cards? Why is only one normal distribution table needed to find any probability under the normal curve? What are chances of getting heads 24 times in a row? How would you put probably in a sentence? What is probability of a drawn card being a club? How many ways can you roll number 7 with 2 dice?